Brookline Specialty Hospital, a not-for-profit provider, had revenues of $12 million in 2018. Expenses other than depreciation totaled 75% of revenues. Brookline purchased fixed assets for $50 million in 2010 and expects to sell these assets for $20 million at the end of their 20-year lifetime.

a. Calculate Brookline’s annual depreciation expense.

b. Construct Brookline’s income statement for Year Ended December 31, 2018.

c. Calculate Brookline’s net income, total profit margin, and approximate cash flow. In one sentence, provide an economic interpretation of its total profit margin.

d. Assume that Brookline is instead organized as a for-profit corporation. It has the same situation as reported in the original problem, but it now must pay taxes at a rate of 40% of pre-tax income. Calculate Brookline’s net income, total profit margin, and approximate cash flow.
Answer #1

Given,

Revenues = $12 million

Expenses other than depreciation = 75% of revenues = 75% * $12 million = $9 million

Expenses other than depreciation = $9 million

Question (a):

Purchase price of fixed assets in 2010 = $50 million

Residual value (re-sale value) at the end of 20th year = $20 million

Therefore, useful life of the fixed assets = 20 years (from 2010)

Depreciation = Purchase Price - Residual value Life.of.the. Asset

50million - 20million Depreciation = 20

30million Depreciation = 20

Depreciation = 1.5million

Therefore, Annual Depreciation Expense = $1.5 million.

Question (b):

Brookline's Income Statement for the Year Ended December 31, 2018
Particulars Amount ($ million) Particulars Amount ($ million)
Expenses other than depreciation 9 Revenues/ Sales 12
Depreciation 1.5
Net income
(12-9-1.5)
1.5
Total 12 Total 12

Question (c):

From the above income statement, Brookline's Net income = $1.5 million

Total profit margin:

Total profit means the net revenue that remains once all costs have been deducted.

Therefore, total profit = net income = $1.5 million

Net Income Total ProfitMargin = * 100 Sales

Total ProfitMargin = 1.5million 12million -* 100

Total ProfitMargin = 0.125 * 100

Therefore, total profit margin = 12.5%

Approximate cash flow = net income + non cash expenses

Here, depreciation is the only non cash expense we have.

Approximate cash flow = $1.5 million + $1.5 million

Approximate cash flow = $3 million

Being a non-profit corporation, it is generating a 12.5% profit. This can be used for development purpose of the organisation and also can be used for helping others by coming up with new plans.

Question (d):

Assuming that Brookline is a for-profit organisation.

Given tax rate = 40%

pre-taxed income = $1.5 million

Tax = 40% of pre-tax income = 40% * $1.5 million = $0.6 million

After-tax income = Pre-tax income - tax = $1.5 million - $0.6 million = $0.9 million

Therefore net income = $0.9 million

Total profit margin = [ $0.9 million/ $12 million ] * 100

Total profit margin = 0.075 * 100 = 7.5%

Total profit margin = 7.5%

Cash flow = net income + depreciation = $0.9 million + $1.5 million = $2.4 million

Approximate cash flow = $2.4 million